# Humanoid Policy $\sim$ Human Policy

### v1

@@ Abstract
Training manipulation policies for humanoid robots with diverse data enhance their robustness and generalization across tasks and platforms. However, learning solely from robot demonstrations is labor-intensive, requiring expensive tele-operated data collection which is difficult to scale. This paper investigates a more scalable data source, egocentric human demonstrations, to serve as cross-embodiment training data for robot learning. We mitigate the embodiment gap between humanoids and humans from both the data and modeling perspectives. We collect an egocentric task-oriented dataset (PH2D) that is directly aligned with humanoid manipulation demonstrations. We then train a human-humanoid behavior policy, which we term Human Action Transformer (HAT). The state-action space of HAT is unified for both humans and humanoid robots and can be differentiably retargeted to robot actions. Co-trained with smaller-scale robot data, HAT directly models humanoid robots and humans as different embodiments without additional supervision. We show that human data improves both generalization and robustness of HAT with significantly better data collection efficiency.
@@ I Introduction
In this paper, we introduce to perform cross-human and humanoid embodiment training for robotic manipulation. Our key insight is to model bimanual humanoid behaviors by directly imitating human behaviors without using learning surrogates such as affordances [40, 3]. To realize this, we first collect an egocentric task-oriented dataset of Physical Humanoid-Human Data, dubbed PH2D. We adapt consumer-grade VR devices to collect egocentric videos with automatic but accurate hand pose and end effector (i.e., hand) annotations. Compared to existing human daily behavior datasets [22, 12], PH2D is task-oriented so that it can be directly used for co-training. The same VR hardwares are then used to perform teleoperation to collect smaller-scale humanoid data for better alignment. We then train a Human-humanoid Action Transformer (HAT), which predicts future hand-finger trajectories in a unified human-centric state-action representation space. To obtain robot actions, we simply apply inverse kinematics and hand retargeting to convert human actions to robot actions. Such a conversion is a differentiable process that allows end-to-end training on different embodiments.
@@ II Related Work
Imitation Learning for Robot Manipulation. Recently, learning robot policy with data gathered directly from the multiple and target robot embodiment has shown impressive robustness and dexterity [71, 43, 56, 37, 10, 49, 9, 38] compared to RL- and planning-based approaches [48, 28, 36, 35, 68] due to direct alignment with the real world. The scale of data for imitation learning has grown substantially with recent advancements in data collection [1, 9, 11, 62], where human operators can efficiently collect large amounts of high-quality, task-oriented data. Despite these advances, achieving open-world generalization still remains a significant challenge due to lack of internet-scale training data.
@@ III-A PH2D: Task-oriented Physical Humanoid-Human Data
We collect task-oriented egocentric human demonstrations by asking human operators to perform tasks overlapping with robot execution (e.g., grasping and pouring) when wearing the VR devices. For every demonstration, we provide language instructions (e.g., grasp a can of coke zero with right hand), and synchronize proprioception inputs and visual inputs by closet timestamps.
@@ III-B HAT: Human Action Transformer
where $\lambda=2$ is an (insensitive) hyperparameter used to balance loss to emphasize the importance of end effector positions over learning unnecessarily precise finger tip keypoints.
@@ IV-B Cross-embodiment Generalization
TABLE V: Ablation of how human whole-body movement in training demonstrations affects policy rollour. We collect same number of demonstrations on the same set of objects for the grasping task with or without whole-body movement. Since the robot does not have natural whole-body movement like humans, it negatively influences the manipulation success rate.
@@ IV-C Ablation Study
Normalization of different embodiments. Tab. II suggests minor differences between using different normalization coefficients for the states and actions vectors humans and humanoids. We take a closer look in Fig. 5, where we investigate the impact of different normalization strategies in the vertical grasping (picking) task. Noticeably, the same normalization approach achieved the highest overall success rate, but the success distribution is biased towards the upper-right region of the grid.

### v2

@@ Abstract
Training manipulation policies for humanoid robots with diverse data enhances their robustness and generalization across tasks and platforms. However, learning solely from robot demonstrations is labor-intensive, requiring expensive tele-operated data collection which is difficult to scale. This paper investigates a more scalable data source, egocentric human demonstrations, to serve as cross-embodiment training data for robot learning. We mitigate the embodiment gap between humanoids and humans from both the data and modeling perspectives. We collect an egocentric task-oriented dataset (PH2D) that is directly aligned with humanoid manipulation demonstrations. We then train a human-humanoid behavior policy, which we term Human Action Transformer (HAT). The state-action space of HAT is unified for both humans and humanoid robots and can be differentiably retargeted to robot actions. Co-trained with smaller-scale robot data, HAT directly models humanoid robots and humans as different embodiments without additional supervision. We show that human data improves both generalization and robustness of HAT with significantly better data collection efficiency.
@@ I Introduction
This paper performs cross-human and humanoid embodiment training for robotic manipulation. Our key insight is to model bimanual humanoid behaviors by directly imitating human behaviors without using learning surrogates such as affordances [37, 3]. To realize this, we first collect an egocentric task-oriented dataset of Physical Humanoid-Human Data, dubbed PH2D. We adapt consumer-grade VR devices to collect egocentric videos with automatic but accurate hand pose and end effector (i.e., hand) annotations. Compared to existing human daily behavior datasets [22, 12], PH2D is task-oriented so that it can be directly used for co-training. The same VR hardwares are then used to perform teleoperation to collect smaller-scale humanoid data for better alignment. We then train a Human-humanoid Action Transformer (HAT), which predicts future hand-finger trajectories in a unified human-centric state-action representation space. To obtain robot actions, we simply apply inverse kinematics and hand retargeting to convert human actions to robot actions. Such a conversion is a differentiable process that allows end-to-end training on different embodiments.
@@ II Related Work
Imitation Learning for Robot Manipulation. Recently, learning robot policy with data gathered directly from the multiple and target robot embodiment has shown impressive robustness and dexterity [66, 40, 52, 34, 10, 45, 9, 35]. The scale of data for imitation learning has grown substantially with recent advancements in data collection [1, 9, 11, 58], where human operators can efficiently collect large amounts of high-quality, task-oriented data. Despite these advances, achieving open-world generalization still remains a significant challenge due to lack of internet-scale training data.
@@ III-A PH2D: Task-oriented Physical Humanoid-Human Data
We collect task-oriented egocentric human demonstrations by asking human operators to perform tasks overlapping with robot execution (e.g., grasping and pouring) when wearing the VR devices. For every demonstration, we provide language instructions (e.g., grasp a can of coke zero with right hand), and synchronize proprioception inputs and visual inputs by closest timestamps.
@@ III-B HAT: Human Action Transformer
where $\pi(s_{i})_{\text{EEF}}$ and $a_{i,\text{EEF}}$ are predicted 6dof end-effector poses to emphasize the importance of end effector positions over learning unnecessarily precise finger tip keypoints, $\lambda=2$ is an (insensitive) hyperparameter used to balance loss.
@@ IV-B Cross-embodiment Generalization
TABLE V: Ablation of how human whole-body movement in training demonstrations affects policy rollout. We collect the same number of demonstrations on the same set of objects for the grasping task with or without whole-body movement. Since the robot does not have a natural whole-body movement like humans, it negatively influences the manipulation success rate.
@@ IV-C Ablation Study
Normalization of different embodiments. Tab. II suggests minor differences between using different normalizations of states and actions for humans and humanoids. We take a closer look in Fig. 5, where we investigate the impact of different normalization strategies in the vertical grasping (picking) task. Noticeably, the same normalization approach achieved the highest overall success rate, but the success distribution is biased towards the upper-right region of the grid.